Problem 11A-8 Applying Overhead; Overhead Variances [LO11-3, LO11-4] Lane Company manufactures a single product that requires a great deal of hand labor. Overhead cost is applied on the basis of standard direct labor-hours. Variable manufacturing overhead should be $2.40 per standard direct labor-hour and fixed manufacturing overhead should be $384,000 per year. The company's product requires 4 pounds of material that has a standard cost of $4.00 per pound and 1.5 hours of direct labor time that has a standard rate of $12.20 per hour. The company planned to operate at a denominator activity level of 60,000 direct labor-hours and to produce 40,000 units of product during the most recent year. Actual activity and costs for the year were as follows Number of units produced Actual direct labor-hours worked Actual variable manufacturing overhead cost incurred Actual fixed manufacturing overhead cost incurred 48,000 78,000 $124,800 $429,000 Required: 1. Compute the predetermined overhead rate for the year. Break the rate down into variable and fixed elements. (Round your answers to 2 decimal places.) Predetermined overhead rate Variable rate $8.80 per DLH $2.40 per DLH $6.40 per DLH Fixed rate
2. Prepare a standard cost card for the company's product. (Round your answers to 2 decimal places.) Direct materials Direct labor Variable overhead Fixed overhead Standard cost per unit 4.00|pounds at 「$ 4.00|per pound | $ 16.00 18.30 3.60 9.60 $ 47.50 1.50 DLHs at 1.50|DLHs at 1.50|DLHs at 「$ 「$ $ 12.20 per DLH 2.40|per DL 6.40 per DLH 3a. Compute the standard direct labor-hours allowed for the year's production Standard direct labor hours72,000 3b. Complete the following Manufacturing Overhead T-account for the year: Manufacturing Overhead 124,800 429,000 79,800 633,600
4. Determine the reason for the underapplied or overapplied overhead from (3) above by computing the variable overhead rate and efficiency variances and the fixed overhead budget and volume variances. (Indicate the effect of each variance by selecting "F" for favorable, "Ufor unfavorable, and "None" for no effect (i.e., zero variance).) Variable overhead rate variance Variable overhead efficiency variance Fixed overhead budget variance Fixed overhead volume variance
Answer #1

Solution 1:

Computation of Predetermined overhead rate - Lane Company
Particulars Amount
Budgeted variable manufacturing overhead ($2.40*60000) $1,44,000
Budgeted fixed manufacturing overhead $3,84,000
Total Manufacturing overhead $5,28,000
/Budgeted direct labor hours 60000
Predetermined overhead rate $8.80
Less: Variable overhead rate (Per direct labor hour) $2.40
Fixed manufacturing overhead rate (per direct labor hour) $6.40

Solution 2:

Standard Cost Card - Lane Company
Direct Material 4 Pounds at $4.00 per pound $16.00
Direct labor 1.5 DLHs at $12.20 per DLH $18.30
Variable manufacturing overhead 1.5 DLHs at $2.40 per DLH $3.60
Fixed manufacturing overhead 1.5 DLHs at $6.40 per DLH $9.60
Standard cost per unit $47.50

Solution 3-a:

Standard Hours Allowed = 48000*1.5 = 72000 hours

Solution 3-b:

Manufacturing Overhead
Particulars Debit Particulars Credit
To Cash (Variable manufacturing overhead) $1,24,800 By WIP (Applied overhead) (72000*$8.80) $6,33,600
To Cash (Fixed manufacturing overhead) $4,29,000
To overhead variance (Overapplied overhead) $79,800
Total $6,33,600 Total $6,33,600

Solution 4:

Variable overhead actual rate = $124800/ 78000 = $1.60 per hour

Variable overhead rate variance = (SP - AP) *Actual Hours = ($2.40 - $1.60) * 78000 = $62,400 Favorable

Variable overhead Efficiency variance = (Standard hours - Actual Hours) *SP = (72000 - 78000)*$2.40 = - $14,400 (Unfavorable)

Fixed overhead budget Variance = Budgeted Fixed Overhead - Actual Fixed overhead = $384000 - $429000 = -$45,000 (unfavorable)

Fixed Overhead Volume Variance = ($6.40*72000) - $384000 = $76,800 Favorable